Circular No. 6276

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GALACTIC CENTER J. D. Kurfess and J. E. Grove, Naval Research Laboratory; D. Messina,

SFA, Inc.; and J. Tueller, Goddard Space Flight Center, report: “We have acquired further Compton Observatory observations that further constrain the position of the source first reported on IAUC 6272. Data obtained during Dec. 15–17 with the four OSSE detectors scanning along the galactic plane provide a galactic-longitude position between −0^{o}.27 and +0^{o}.17 (95percent confidence). The OSSE field-of-view normal to the scan direction is 11^{o}.4 FWHM. When coupled with the position reported on IAUC 6275, the location of the transient repeater is constrained to a region with the following vertices (equinox 2000.0): α = 17^{h}44^{m}34^{s}, δ = −28^{o}34′.6; 17^{h}45^{m}58^{s}, −28^{o}46′.4; 17^{h}43^{m}11^{s}, −29^{o}01′.6; 17^{h}41^{m}43^{s}, −28^{o}49′.0.”

COMET C/1995 O1 (HALE-BOPP) M. Womack, Pennsylvania State University at Erie; S. A. Stern, South-

west Research Institute; and M. C. Festou, Observatoire Midi Pyrenees, report observations with the National Radio Astronomy Observatory 12-m telescope at Kitt Peak: “The CO J=2–1 line at 230 GHz was observed in this comet (at r = 5.9 AU) on Dec. 8, 9, and 10 UT. The J=2–1 line was not detected on Dec. 8, down to a 3σ limit of T dv = 0.049 K km/s, but it was detected with integrated line flux areas of T dv = 0.028 ± 0.004 and 0.082 ± 0.004 K km/s, on Dec. 9 and 10, respectively. The T dv for Dec. 9 is a factor of 2–3 times lower than reported for this comet in August and September (IAUC 6234, 6236). The data correspond to production rates of Q(CO) ∼ 5 and 15 × 10^{27} mol s^{−1} for Dec. 9 and 10, respectively. During the last few hours of Dec. 10, the emission exhibited an increase in line intensity of a factor of 4, but the integrated flux was conserved. The average linewidth on Dec. 9 and 10 is 0.4 km/s and the line appears blue-shifted by ∼ 0.36 km/s.”

COMET 73P/SCHWASSMANN-WACHMANN 3 Total visual magnitude estimates (cf. IAUC 6259): Nov. 24.10 UT,

7.4 (A. Hale, Cloudcroft, NM, 10×50 binoculars); Dec. 16.12, 8.2 (C. S. Morris, Whitaker Peak, CA, 20×80 binoculars); 21.07, 7.8 (R. Keen, Mt. Thorodin, CO, 7×50 binoculars).

1995 December 21 Daniel W. E. Green